Circular No. 8190

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN SCUTUM S. Nakano, Sumoto, Japan; and H. Sato, National Astronomical Obser-

vatory of Japan (NAOJ), report the discovery of a possible nova by Hideo Nishimura (Kakegawa, Shizuoka-ken, Japan) on T-Max films taken on Aug. 28.58 (mag 8.5) and 29.436 UT (mag ∼ 8.4) with a Pentax 200-mm f/4.0 lens, the position given as α = 18^{h}49^{m}38^{s}, δ = −9^{o}33′45′′ (equinox 2000.0). T. Nakamura, NAOJ, reports the following mean precise position end figures by S. Wakuda (Hamamatsu, Japan) from six CCD images (0.25-m f/5.1 reflector, Aug. 29.354, V = 9.5): 37^{s}.62, 50′′.3. H. Yamaoka, Kyushu University, reports that ASAS-3 images show an object at position end figures 37^{s}.7, 53′′ with the following V magnitudes: Aug. 24.105, [14.0; 26.096, 10.15; 28.096, 8.94. Yamaoka adds that a USNO-A2.0 star of red mag 16.4 and blue mag 17.2 has position end figures 37^{s}.684, 53′′.76; this star is also present in the GSC-2.2 and USNO-B1.0 catalogues, and a 2MASS star (J = 14.32, K = 13.82) has position end figures 37^{s}.70, 54′′.1. Visual magnitude estimate by A. Pearce, Nedlands, W. Australia: Aug. 30.461, 9.3.

CXOU J180951.1–194351 = RXTE J1810–197 E. V. Gotthelf and J. P. Halpern, Columbia University; C. Markwardt,

Goddard Space Flight Center (GSFC); A. I. Ibrahim, George Washington University and GSFC; M. Roberts and S. M. Ransom, McGill University; and P. Woods, USRA and National Space Science and Technology Center, on behalf of a larger collaboration, report a highly significant detection of a pulsed source with a barycentric period of 5.5391(3) s, from a 2833-s Chandra HRC-I observation starting Aug. 27.95166 UT. This source shows a strongly modulated (52-percent pulsed fraction) sinusoidal pulse signal with a countrate of 0.917 counts/s in the HRC-I energy band 0.08–10.0 keV. The pulsar is located at α = 18^{h}09^{m}51^{s}.13, δ = −19^{o}43′51′′.7 (equinox 2000.0; uncertainty radius of 2′′.1 at the off-axis HRC-I position). This position is offset by 14′ from the centroid reported for the 5.5-s pulsar RXTE J1810–197 (IAUC 8168); the large offset (6σ) is likely due to systematic errors in the derived RXTE position which are based on scans of the complex region. The location of CXOU J180951.1–194351 is consistent with a source seen in archival ROSAT and ASCA images. The lack of an optical counterpart within the 1σ Chandra error circle suggests that CXOU J180951.1–194351 is an isolated neutron star, one most similar to the transient anomalous x-ray pulsar AX J1844.8–0256 (cf. Vasisht et al. 2000, Ap.J. 43, L49).

2003 August 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green